OpenStudy (anonymous):

Calculate the area of the regular pentagon below: A regular pentagon with side length of 22.3 inches and dotted line from center to middle of side of 15.4 inches. 686.84 square inches 732.34 square inches 858.55 square inches 1717.1 square inches @StudyGurl14 @gold_grace123

OpenStudy (studygurl14):

Pentagon - five sides legnth of one side - 22.3 apothem: 15.4 Area of pentagon: \(\large\frac{perimeter\times apothem}{2}\)

OpenStudy (anonymous):

huh @StudyGurl14

OpenStudy (studygurl14):

Do this: \(\large\frac{5(22.3)\times15.4}{2}\)

OpenStudy (anonymous):

i dont get it still @StudyGurl14

OpenStudy (studygurl14):

serisouly? Just multiply 5 and 22.3 and 15.4. Then divide that by 2. Boom that's your answer